Public television station KQED in San Francisco broadcasts a sinusoidal radio signal at a power of 316 kW. Assume that the wave spreads out uniformly into a hemisphere above the ground.

A) At a home 5.50km away from the antenna, what average pressure does this wave exert on a totally reflecting surface? answer in PA units

B)At a home 5.50km away from the antenna, what are the amplitudes of the electric and magnetic fields of the wave? answer in N/C, T

C) At a home 5.50km away from the antenna, what is the average density of the energy this wave carries? answer in J/m^3

D) For the energy density in part (c), what percentage is due to the electric field?

E)For the energy density in part (c), what percentage is due to the magnetic field?
Answer #1
Concepts and reason

The concepts required to solve this problem is properties of electromagnetic waves.

Initially, calculate the average pressure by using the intensity of the wave. Later, use the amplitude equation to calculate the electric and magnetic field amplitude. Then, calculate the average density of the energy of the electromagnetic wave. Finally, calculate the energy density due to electric and magnetic field and their percentage in the total energy density of the wave.

Fundamentals

The intensity of electromagnetic wave is given as,

I=PAI = \frac{P}{A}

Here, P is the power, and AA is the surface area over which wave spreads.

The amplitude of electric field is given as,

Emax=2Iε0c{E_{\max }} = \sqrt {\frac{{2I}}{{{\varepsilon _0}c}}}

Here, ε0{\varepsilon _0} is the vacuum permittivity, II is the intensity of the incident light, and cc is the speed of light.

The average energy density of the electromagnetic wave is,

u=12ε0Eavg2+12μ0Bavg2u = \frac{1}{2}{\varepsilon _0}E_{{\rm{avg}}}^2 + \frac{1}{{2{\mu _0}}}B_{{\rm{avg}}}^2

Here, ε0{\varepsilon _0} is the vacuum permittivity, μ0{\mu _0} is the vacuum permeability, Eavg{E_{{\rm{avg}}}} is the average value of the electric field, and Bavg{B_{{\rm{avg}}}} is the average value of the magnetic field.

The energy density due to electric field is,

uE=12ε0Eavg2{u_{\rm{E}}} = \frac{1}{2}{\varepsilon _0}E_{{\rm{avg}}}^2

Here, ε0{\varepsilon _0} is the vacuum permittivity and Eavg{E_{{\rm{avg}}}} is the average value of the electric field.

The energy density due to magnetic field is,

uB=12μ0Bavg2{u_{\rm{B}}} = \frac{1}{{2{\mu _0}}}B_{{\rm{avg}}}^2

Here, μ0{\mu _0} is the vacuum permeability and Bavg{B_{{\rm{avg}}}} is the average value of the magnetic field.

(A)

Substitute 2πr22\pi {r^2} for AA in the equation I=PAI = \frac{P}{A} .

I=P2πr2I = \frac{P}{{2\pi {r^2}}}

Substitute 316kW316{\rm{ kW}} for PP and 5.50km5.50{\rm{ km}} for rr in the equation I=P2πr2I = \frac{P}{{2\pi {r^2}}} and calculate the intensity of the electromagnetic wave.

I=(316kW)2π(5.50km)2=(316kW(103W1kW))2π(5.50km(103m1km))2=1.66×103W/m2\begin{array}{c}\\I = \frac{{\left( {316{\rm{ kW}}} \right)}}{{2\pi {{\left( {5.50{\rm{ km}}} \right)}^2}}}\\\\ = \frac{{\left( {316{\rm{ kW}}\left( {\frac{{{{10}^3}{\rm{ W}}}}{{1{\rm{ kW}}}}} \right)} \right)}}{{2\pi {{\left( {5.50{\rm{ km}}\left( {\frac{{{{10}^3}{\rm{ m}}}}{{1{\rm{ km}}}}} \right)} \right)}^2}}}\\\\ = 1.66 \times {10^{ - 3}}{\rm{ W/}}{{\rm{m}}^2}\\\end{array}

Use the pressure equation.

Substitute 1.66×103W/m21.66 \times {10^{ - 3}}{\rm{ W/}}{{\rm{m}}^2} for II and 3.0×108m/s3.0 \times {10^8}{\rm{ m/s}} for cc in the equation p=2Icp = \frac{{2I}}{c} and calculate the average pressure due to electromagnetic wave for a totally reflecting surface.

p=2(1.66×103W/m2)3.0×108m/s=1.1×1011pa\begin{array}{c}\\p = \frac{{2\left( {1.66 \times {{10}^{ - 3}}{\rm{ W/}}{{\rm{m}}^2}} \right)}}{{3.0 \times {{10}^8}{\rm{ m/s}}}}\\\\ = 1.1 \times {10^{ - 11}}{\rm{ pa}}\\\end{array}

(B)

Use the electric field amplitude equation.

Substitute 1.66×103W/m21.66 \times {10^{ - 3}}{\rm{ W/}}{{\rm{m}}^2} for II , 8.854×1012C2/Nm28.854 \times {10^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2} for ε0{\varepsilon _0} , and 3.0×108m/s3.0 \times {10^8}{\rm{ m/s}} for cc in the equation Emax=2Iε0c{E_{\max }} = \sqrt {\frac{{2I}}{{{\varepsilon _0}c}}} .

Emax=2(1.66×103W/m2)(8.854×1012C2/Nm2)(3.0×108m/s)=1.12N/C\begin{array}{c}\\{E_{\max }} = \sqrt {\frac{{2\left( {1.66 \times {{10}^{ - 3}}{\rm{ W/}}{{\rm{m}}^2}} \right)}}{{\left( {8.854 \times {{10}^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2}} \right)\left( {3.0 \times {{10}^8}{\rm{ m/s}}} \right)}}} \\\\ = 1.12{\rm{ N/C}}\\\end{array}

The amplitude of the magnetic field is,

Bmax=Emaxc{B_{\max }} = \frac{{{E_{\max }}}}{c}

Substitute 1.12N/C1.12{\rm{ N/C}} for Emax{E_{\max }} and 3.0×108m/s3.0 \times {10^8}{\rm{ m/s}} for cc in the equation Bmax=Emaxc{B_{\max }} = \frac{{{E_{\max }}}}{c} .

Bmax=1.12N/C3.0×108m/s=3.73×109T\begin{array}{c}\\{B_{\max }} = \frac{{1.12{\rm{ N/C}}}}{{3.0 \times {{10}^8}{\rm{ m/s}}}}\\\\ = 3.73 \times {10^{ - 9}}{\rm{ T}}\\\end{array}

(C)

Use the average energy density equation

Substitute Emax2\frac{{{E_{\max }}}}{{\sqrt 2 }} for Eavg{E_{{\rm{avg}}}} and Bmax2\frac{{{B_{\max }}}}{{\sqrt 2 }} for Bavg{B_{{\rm{avg}}}} in the equation u=12ε0Eavg2+12μ0Bavg2u = \frac{1}{2}{\varepsilon _0}E_{{\rm{avg}}}^2 + \frac{1}{{2{\mu _0}}}B_{{\rm{avg}}}^2 .

u=14ε0Emax2+14μ0Bmax2u = \frac{1}{4}{\varepsilon _0}E_{\max }^2 + \frac{1}{{4{\mu _0}}}B_{\max }^2

Substitute 8.854×1012C2/Nm28.854 \times {10^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2} for ε0{\varepsilon _0} , 1.12N/C1.12{\rm{ N/C}} for Emax{E_{\max }} , 4π×107Tm/A4\pi \times {10^{ - 7}}{\rm{ T}} \cdot {\rm{m/A}} for μ0{\mu _0} , and 3.73×109T3.73 \times {10^{ - 9}}{\rm{ T}} for Bmax{B_{\max }} in the equation u=14ε0Emax2+14μ0Bmax2u = \frac{1}{4}{\varepsilon _0}E_{\max }^2 + \frac{1}{{4{\mu _0}}}B_{\max }^2 .

u=14(8.854×1012C2/Nm2)(1.12N/C)2+14(4π×107Tm/A)(3.73×109T)2=5.54×1012J/m3\begin{array}{c}\\u = \frac{1}{4}\left( {8.854 \times {{10}^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2}} \right){\left( {1.12{\rm{ N/C}}} \right)^2} + \frac{1}{{4\left( {4\pi \times {{10}^{ - 7}}{\rm{ T}} \cdot {\rm{m/A}}} \right)}}{\left( {3.73 \times {{10}^{ - 9}}{\rm{ T}}} \right)^2}\\\\ = 5.54 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}\\\end{array}

(D)

Use the average energy density equation

Substitute Emax2\frac{{{E_{\max }}}}{{\sqrt 2 }} for Eavg{E_{{\rm{avg}}}} in the equation uE=12ε0Eavg2{u_{\rm{E}}} = \frac{1}{2}{\varepsilon _0}E_{{\rm{avg}}}^2 .

uE=14ε0Emax2{u_{\rm{E}}} = \frac{1}{4}{\varepsilon _0}E_{\max }^2

Substitute 8.854×1012C2/Nm28.854 \times {10^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2} for ε0{\varepsilon _0} , and 1.12N/C1.12{\rm{ N/C}} for Emax{E_{\max }} in the equation uE=14ε0Emax2{u_{\rm{E}}} = \frac{1}{4}{\varepsilon _0}E_{\max }^2 .

uE=14(8.854×1012C2/Nm2)(1.12N/C)2=2.77×1012J/m3\begin{array}{c}\\{u_{\rm{E}}} = \frac{1}{4}\left( {8.854 \times {{10}^{ - 12}}{\rm{ }}{{\rm{C}}^2}{\rm{/N}} \cdot {{\rm{m}}^2}} \right){\left( {1.12{\rm{ N/C}}} \right)^2}\\\\ = 2.77 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}\\\end{array}

The percentage of the energy due to electric field in the total energy density is,

percentage=uEu×100%percentage = \frac{{{u_{\rm{E}}}}}{u} \times 100\%

Substitute 2.77×1012J/m32.77 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3} for uE{u_{\rm{E}}} and 5.54×1012J/m35.54 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3} for uu in the equation percentage=uEu×100%percentage = \frac{{{u_{\rm{E}}}}}{u} \times 100\% .

percentage=2.77×1012J/m35.54×1012J/m3×100%=50%\begin{array}{c}\\percentage = \frac{{2.77 \times {{10}^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}}}{{5.54 \times {{10}^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}}} \times 100\% \\\\ = 50\% \\\end{array}

(E)

Use the average energy density equation

Substitute Bmax2\frac{{{B_{\max }}}}{{\sqrt 2 }} for Bavg{B_{{\rm{avg}}}} in the equation uB=12μ0Bavg2{u_{\rm{B}}} = \frac{1}{{2{\mu _0}}}B_{{\rm{avg}}}^2 .

uB=14μ0Bmax2{u_{\rm{B}}} = \frac{1}{{4{\mu _0}}}B_{\max }^2

Substitute 4π×107Tm/A4\pi \times {10^{ - 7}}{\rm{ T}} \cdot {\rm{m/A}} for μ0{\mu _0} , and 3.73×109T3.73 \times {10^{ - 9}}{\rm{ T}} for Bmax{B_{\max }} in the equation uB=14μ0Bmax2{u_{\rm{B}}} = \frac{1}{{4{\mu _0}}}B_{\max }^2 .

uB=14(4π×107Tm/A)(3.73×109T)2=2.77×1012J/m3\begin{array}{c}\\{u_{\rm{B}}} = \frac{1}{{4\left( {4\pi \times {{10}^{ - 7}}{\rm{ T}} \cdot {\rm{m/A}}} \right)}}{\left( {3.73 \times {{10}^{ - 9}}{\rm{ T}}} \right)^2}\\\\ = 2.77 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}\\\end{array}

The percentage of the energy due to magnetic field in the total energy density is,

percentage=uBu×100%percentage = \frac{{{u_{\rm{B}}}}}{u} \times 100\%

Substitute 2.77×1012J/m32.77 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3} for uB{u_{\rm{B}}} and 5.54×1012J/m35.54 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3} for uu in the above equation.

percentage=2.77×1012J/m35.54×1012J/m3×100%=50%\begin{array}{c}\\percentage = \frac{{2.77 \times {{10}^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}}}{{5.54 \times {{10}^{ - 12}}{\rm{ J/}}{{\rm{m}}^3}}} \times 100\% \\\\ = 50\% \\\end{array}

Ans: Part A

The average pressure due to incident intensity of electromagnetic wave for a totally reflecting surface is 1.1×1011pa1.1 \times {10^{ - 11}}{\rm{ pa}} .

Part B

The amplitude of electric field is 1.12N/C1.12{\rm{ N/C}} and magnetic field is 3.73×109T3.73 \times {10^{ - 9}}{\rm{ T}} .

Part C

The average density of the energy this wave carries is 5.54×1012J/m35.54 \times {10^{ - 12}}{\rm{ J/}}{{\rm{m}}^3} .

Part D

The percentage of energy density due to the electric field is 50%.

Part E

The percentage of energy density due to the magnetic field is 50%.